# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v1

@@ 3.2. Adds redundancy in the serving framework
PagedAttention makes an LLM serving system responsible for managing the mappings between KV-cache and dynamically allocated memory blocks. For example, consider a request that allocates four KV-cache blocks over time (left half of Figure 2). These blocks are usually non-contiguous in virtual memory. During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV-cache blocks. To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime. This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half Figure 2).
@@ 3.3. Performance Overhead
PagedAttention also leads to potential performance issues both on GPU and CPU. We investigate them separately.
@@ 3.3.1. Runtime overhead on the GPU
PagedAttention slows down attention computation by adding extra code in the critical path. For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (Kwon et al., 2023). Figure 3 shows that the paged decode kernel in FlashAttention is also slower than the vanilla kernel. Our further analysis reveals that the number of instructions executed in PagedAttention is up to $13\%$ higher than the vanilla kernel. We also find that the overhead of paging reduces at high batch sizes or long context lengths. This is because computing attention for decodes is memory bound and when the KV-cache size is large, memory stalls hide the instruction overhead.
Figure 4. Overhead of PagedAttention in FlashInfer’s prefill kernel (model: Yi-6B). Legend shows the batch size.
However, hiding instruction overhead is more challenging in a prefill attention kernel which is compute-bound with $N^{2}$ complexity. For example, Figure 4 shows that the paged version of FlashInfer (Ye et al., 2024) prefill kernel is up to 14% slower than the vanilla kernel. The paged kernel is also missing some well-known optimizations (fip, 2024).
To highlight another example of difficulty involved in writing an efficient attention kernel, Figure 5 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128. Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
@@ 3.3.2. Runtime overhead on the CPU
To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel. In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV-cache blocks in the longest request of the batch. This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV-cache blocks in each request by padding unoccupied slots with zeros. If a batch contains a few long and many short requests, such padding results in a significant overhead. In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations. While a recent fix (vll, 2024a) has mitigated some of this overhead, we find that it can still be as high as $10\%$. High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (trt, 2023). This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead.
Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient. vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging. However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads in terms of memory management.
Figure 5. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Yi-6B). Legend shows the block size.
(a) Output tokens per second.
(b) Decode iteration latency.
(c) Memory allocation bandwidth.
Figure 6. LLM inference throughput plateaus at large batch sizes (a) beyond which latency suffers (b) without improving throughput. This also means that the rate at which physical memory needs to be allocated also saturates (c).
@@ 4. Insights into LLM Serving Systems
To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single NVIDIA A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism. We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure latency, throughput and memory requirement of the decode phase (see §6 for our discussion and optimizations for the prefill phase).
Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance. This is due to auto-regressive decoding that generates one token at a time, per-request. Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token until the request completes.
Observation-2: KV-cache does not require high memory allocation bandwidth. The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory. For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively. Further, each iteration runs for 10s-100s of milliseconds (6(b)) implying that a request requires at most a few megabytes of memory per second. While batching improves system throughput, the number of tokens generated per second plateaus beyond a certain batch size (6(a)). This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B). For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (6(c)).
In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache. In the next section, we begin with a high-level design description of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§6).
5. vAttention: System Design
Our goal is to improve efficiency and portability by adding dynamic memory allocation support to existing kernels. To achieve this goal, vAttention leverages system support for dynamic memory allocation instead of implementing paging in user space.
@@ 5.1. Design Overview
vAttention builds on the ability to allocate virtual memory and physical memory separately. Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime i.e., allocate physical memory only when required (similar to PagedAttention). This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory. This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual address space to each process33364-bits systems typically utilize 48 bits for virtual addresses, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
@@
5.1.1. Pre-reserving virtual memory
Since virtual memory is abundant, we pre-allocate enough virtual memory space that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported.
Number of virtual memory buffers: Each layer in an LLM maintains its own K and V tensors: we refer to them individually as K-cache and V-cache. We allocate separate virtual memory buffers for K-cache and V-cache. For a single GPU job, this requires pre-reserving $2\times N$ buffers where $N$ is the number of layers in the model. In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
Size of a virtual memory buffer: The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker. Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Note that $S$ is constant for a given model configuration.
Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2). In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$. For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$. Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$). Therefore, the total virtual memory requirement for 60 layers of Yi-34B is 120 buffers of 100GB each (12TB total). Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB. Therefore, virtual memory allocations can be satisfied easily.
5.1.2. On-demand physical memory allocation
vAttention preferentially allocates physical memory one page at a time and only when a request has used all of its previously allocated physical memory pages. To show how it works, we refer to a simple example in Figure 7. The example shows how vAttention manages the K-cache (or V-cache) at one layer of the model assuming maximum batch size of two. Rest of the K-cache and V-cache buffers are managed similarly at all layers.
5.2. Leveraging Low-level CUDA Support
The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time. However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cud, 2024; Guo et al., 2024). We leverage these low-level APIs in vAttention.
5.2.1. CUDA virtual memory APIs
Table 3 provides a high-level overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory (see the leftmost column). The allocation granularity depends on the page size used by the GPU and the size of virtual memory buffer or a physical memory handle must be a multiple of the allocation granularity. Different sub-regions of a virtual memory buffer can be backed by physical memory independently of other sub-regions in that buffer (see Figure 7c for an example). For simplicity, we refer to the granularity at which physical memory is allocated as page size.
@@ 5.2.1. CUDA virtual memory APIs
CUDA APIs vAttention APIs Description 64KB 128KB 256KB 2MB
@@
Table 3. Low-level APIs for virtual memory management and their latency with different allocation sizes. * represents APIs that we use once while instantiating or terminating the serving framework. Rest of the APIs are used for (un)mapping physical memory pages at runtime. CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
5.2.2. Extending PyTorch caching allocator:
@@
5.2.3. Request-level KV-cache indexing:
@@ 5.3. Serving LLMs with vAttention
We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers. Our library exposes a set of simple APIs listed in Table 4 to the serving framework.
5.3.1. Initial setup:
@@
Table 4. APIs exposed to a serving framework for dynamic KV-cache memory management with vAttention.
5.3.2. Scheduling a new request:
5.3.3. Model execution:
@@
6. vAttention: Optimizations
@@
There are two primary challenges in using CUDA’s virtual memory support for serving LLMs. First, cuMemCreate currently allocates a minimum of 2MB physical memory page. Large pages can waste physical memory due to internal fragmentation. Second, invoking CUDA APIs incurs high latency. This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
6.1. Mitigating internal fragmentation
@@
We mitigate internal fragmentation by reducing the granularity of physical memory allocation. NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB. Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes. The simplest way to achieve this would be to extend the existing CUDA virtual memory APIs (listed in Table 3) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes). Unfortunately, the CUDA APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
@@
Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source. Therefore, we implement a new set of APIs in the open-source NVIDIA drivers to mimic the same functionality that existing CUDA APIs provide but with support for multiple page sizes. The second column in Table 3 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity. In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes. A serving framework can configure a desired page size in vAttention while initializing it: we recommend using 256KB pages by default. The last set of columns in Table 3 shows the latency of each API with different page sizes.
6.2. Hiding memory allocation latency

### v2

@@ 3.2. Adds Redundancy in the Serving Framework
PagedAttention makes an LLM serving system responsible for managing the mappings between KV-cache and dynamically allocated memory blocks. For example, consider a request that allocates four KV-cache blocks over time (left half of Figure 1). These blocks are usually non-contiguous in virtual memory. During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV-cache blocks. To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime. This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half in Figure 1).
Figure 2. Overhead of PagedAttention in prefill kernels (model: Llama-3-8B, one A100 GPU). Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention (FA) and FlashInfer (FI).
@@ 3.3.1. Runtime overhead on the GPU
PagedAttention slows down attention computation by adding extra code in the critical path. For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (vllmsosp, )).
In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries. For example, PagedAttention based prefill kernels of FlashAttention and FlashInfer are up to 28% and 24% slower than the non-paged kernels in the corresponding libraries. Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels.
To highlight another example of difficulty involved in writing an efficient attention kernel with paging support, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128. Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
@@ 3.3.2. Runtime overhead on the CPU
To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel. In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV-cache blocks in the longest request of the batch. This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV-cache blocks in each request by padding unoccupied slots with zeros. If a batch contains a few long and many short requests, such padding results in a significant overhead. In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations. While a recent fix (vllm-blocktable-issue, ) has mitigated some of this overhead, we find that it can still be as high as $10\%$. High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (tensorrt-perf-decay, ). This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead. However, doing so would be an unwelcome change for most programmers.
Figure 3. Latency of vLLM’s paged decode kernel is sensitive to block size (model: Llama-3-8B, one A100 GPU).
Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient. vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging. However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads from a memory management perspective.
@@ 4. Insights into LLM Serving Systems
To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism. We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §5.4 for our discussion and optimizations for the prefill phase).
Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance. This is a dream property for a memory allocator, thanks to the auto-regressive decoding wherein every forward pass of the model generates one token per request. Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token as long as the request is active.
Observation-2: KV-cache does not require high memory allocation bandwidth. The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory. For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively. Further, each iteration runs for 10s-100s of milliseconds implying that a request requires at most a few megabytes of memory per second. While batching improves system throughput (sarathi2023, ; sarathiserve2024, ; patel2023splitwise, ; distserve2024, ), the number of tokens generated per second plateaus beyond a certain batch size (Figure 4, left). This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B). For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (Figure 4, right).
Figure 4. Decode throughput (left) and the rate of physical memory allocation (right) saturate at large batch sizes.
In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache. In the next section, we begin with a high-level overview of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§5.4).
5. vAttention: Design and Implementation
To facilitate dynamic allocation of physical memory to unmodified attention kernels, vAttention leverages system support for demand paging instead of implementing it in user space.
@@ 5.1. Design Overview
vAttention builds on the ability to allocate virtual memory and physical memory separately. Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime (similar to PagedAttention). This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory. This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual memory per process11164-bits systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
@@
Pre-reserving virtual memory
Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported. In doing so, we assume that each request would generate as many tokens as the maximum sequence length supported by the model.
Number of virtual memory buffers
A serving framework typically maintains separate K and V tensors for each layer of the model: we refer to them individually as K-cache and V-cache. We allocate separate virtual memory buffers for K-cache and V-cache. For a single GPU job, this requires pre-reserving $2\times N$ buffers where $N$ is the number of layers in the model. In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
@@
CUDA VM APIs vAttention VM APIs Description 64KB 128KB 256KB 2MB
@@
Table 2. CUDA VM (virtual memory) APIs. * represents APIs that we use once while instantiating or terminating the serving framework. Rest of the APIs are used for (un)mapping physical memory pages at runtime. CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
Size of a virtual memory buffer
The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker. Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Note that $S$ is constant for a given model configuration.
Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2). In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$. For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$. Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$). Therefore, the total virtual memory requirement for 60 layers is 120 buffers of 100GB each (12TB total). Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB. Therefore, virtual memory allocations can be satisfied easily. Figure 5 shows how vAttention allocates physical memory pages dynamically.
5.2. Leveraging Low-level CUDA Support
The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time. However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cudavirtualmemory, ; gmlakeasplos24, ). We leverage these low-level APIs in vAttention.
5.2.1. CUDA virtual memory APIs
Table 2 provides an overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory. The allocation granularity depends on the page size used by the GPU. Further, the size of a virtual memory buffer or a physical memory handle must be a multiple of the physical memory allocation granularity. It is also important to highlight that physical memory pages can be allocated (or de-allocated) to sub-regions in a virtual memory buffer independently of other sub-regions. For simplicity, we refer to the granularity of physical memory allocation as page size.
5.2.2. Extending PyTorch caching allocator
@@
5.2.3. Request-level KV-cache indexing
@@ 5.3. Serving LLMs with vAttention
We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers. Our library exposes a set of simple APIs to the serving framework (see Table 3 and Algorithm 1).
5.3.1. Initial setup
@@ 5.3.2. Scheduling a new request
Table 3. Key APIs that vAttention exposes to a serving framework for dynamic KV-cache memory management.
5.3.3. Model execution
@@ 5.3.4. Request completion
Algorithm 1 Using vAttention in a serving framework.
1:max_batch_size $\leftarrow$ B
2:cache_seq_len $\leftarrow$ [0]*B
3:req_batch_idx $\leftarrow$ dict()
4:vattention.init_cache(config_params)
5:while !request_pool.is_empty() do
6:     for $R_{i}$ in new_requests do
7:         if can_schedule($R_{i}$then
8:              idx $\leftarrow$ vattention.alloc_reqid()
9:              req_batch_idx[$R_{i}$] $\leftarrow$ idx
10:              cache_seq_len[idx] $\leftarrow$ prompt_len($R_{i}$)
11:         end if
12:     end for
13:     vattention.step(cache_seq_len)
14:     model.forward()
15:     for $R_{i}$ in active_requests do
16:         idx $\leftarrow$ req_batch_idx[$R_{i}$]
17:         if is_complete($R_{i}$then
18:              cache_seq_len[idx] $\leftarrow$ 0
19:              vattention.free_reqid(idx)
20:         else
21:              cache_seq_len[idx] +=+=+ = 1
22:         end if
23:     end for
24:end while
5.4. Optimizations
@@
There are two primary challenges in using CUDA virtual memory support for serving LLMs. First, cuMemCreate currently allocates a minimum of 2MB physical memory page. Large pages can waste physical memory due to internal fragmentation. Second, invoking CUDA APIs incurs high latency. This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
5.4.1. Mitigating internal fragmentation
@@
We mitigate internal fragmentation by reducing the granularity of physical memory allocation. NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (mcmgpus, ; gpu-mismanaged, ; gpu-reverse-engg-tlb1, ; pascalmmu, ). Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes. The simplest way to achieve this would be to extend the existing CUDA virtual memory APIs (listed in Table 2) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes). Unfortunately, the CUDA APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
@@
Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source. Therefore, we implement a new set of APIs in the open-source NVIDIA drivers to mimic the same functionality that existing CUDA APIs provide but with support for multiple page sizes. The second column in Table 2 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity. In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes. A serving framework can configure a desired page size in vAttention while initializing it. The last set of columns in Table 2 shows the latency of each API with different page sizes.
5.4.2. Hiding memory allocation latency